OpenStudy (anonymous):

WILL GIVE MEDAL PLEASE HELP!! :) Which equations model this problem? A recipe that makes 1 serving calls for 0.1 L of honey. How many servings can Alicia make if she has 0.8 L of honey? Use s for the number of servings. Choose all answers that are correct. a. 0.1s=0.8 b. 0.8/ 0.1=s c. 0.8*s=0.1 d. s/ 0.8=0.1

OpenStudy (anonymous):

@texaschic101 :)

OpenStudy (anonymous):

@PaulaLovesSchool13 :)

OpenStudy (anonymous):

thinking o_O #sohard

OpenStudy (anonymous):

haha thx :)

OpenStudy (anonymous):

It's fine if you don't know it @PaulaLovesSchool13 :D ( it's pretty hard)

OpenStudy (anonymous):

o_O I am thinking maybe ask @iGreen

OpenStudy (anonymous):

@iGreen :)

OpenStudy (igreen):

1 serving = 0.1L of Honey We have 0.8L of Honey Therefore we can use the equation: 0.1s = 0.8 to find out how many servings we can make.

OpenStudy (anonymous):

@iGreen lol isn't that the same as b? b. 0.8/ 0.1=s just finding how many 0.1's go into 0.8?

OpenStudy (anonymous):

i was thinking a :/

OpenStudy (anonymous):

Wait he is right, I read the question wrong, they worded it that 0.8 would be the y xD omigosh yeah, just do that.

OpenStudy (anonymous):

They want me to find ALL of the correct anwsers :(

OpenStudy (igreen):

Oh, then it's A and B.

OpenStudy (igreen):

They will both give you the correct answer.

OpenStudy (anonymous):

yeah, because you "A recipe that makes 1 serving calls for 0.1 L of honey. How many servings can Alicia make if she has 0.8 L of honey" she already has 0.8 so just do 0.8/0.1 because 0.1 is the coefficient or in this case the m in the y=mx+b equation.

OpenStudy (anonymous):

oh ok thank you :) I dont mean to bother you but can you help me with a few more

OpenStudy (igreen):

To find out how many servings you must divide \(0.8 \div 0.1\) and Options A and B do that. Option C wants you to divide \(0.1 \div 0.8\) (which is the opposite) Option D wants you to multiply \(0.8 \cdot 0.1\).

OpenStudy (anonymous):

what other questions?

OpenStudy (anonymous):

@iGreen ok

OpenStudy (anonymous):

Annie rappels down a cliff in 3 jumps of 6.5 feet each. Use a positive or negative number to express the change in her altitude.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Answer: She feels hyper

OpenStudy (anonymous):

well you are going down right? so it is just, 6.5 + 6.5 +6.5 or 6.5(3) so that is 19.5, but since you are going down- it's just a negative distance so the change is -19.5

OpenStudy (anonymous):

Incorrect, she feels hyper now

OpenStudy (anonymous):

sorry took a long time, answering something else, but if she was going UP that would be positive, whenever you are going BACKWARDS or DOWN from somewhere, it's minus because you are going back a distance, not gaining.

OpenStudy (anonymous):

yeah hyperactivity can be a converging contingency too xD

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

yeah i guess so lol

OpenStudy (anonymous):

but you get it right?

OpenStudy (anonymous):

i got a 60% :(

OpenStudy (anonymous):

On your test? What questions did you get wrong? :0 :c That's okay, keep trying and ask your professors for extra help- it will get better. Math is about practice. :)